A treasure map required that the seekers, starting from a certain point X, walk 10 m due North, then 17 m due East followed by 25 m due South. Determine the magnitude and the direction of the treasure seekers starting point.


1
Expert's answer
2021-05-24T16:19:46-0400

Gives

Certain point (X)

North direction(AE)= 10 m

East(BC)=17 m

South(CE)=25 m





North to South total distance


Path

DE=CE-AB=25-10=15m

West to east distance =17m

Resultant(AE)= "\sqrt{AD^2+DE^2}"

"AE=\sqrt{17^2+15^2}=\sqrt{514}=22.67m"

Magnitude of direction

AE =22.67m

Resultant direction South -east direction

Where angle "i=\theta"

"tan\theta =\frac{15}{17}"

"\theta=tan^{-1}\frac{15}{17}"

"\theta=41.42°"

Resultant direction South-east

positive x-axis make angle is41.42°